### v2

# Sun-Shine: A Large Language Model for Tibetan Culture

Abstract
Tibetan, a minority language in China, features a highly intricate grammatical structure, characterized by four verb tenses and a tense system with frequent irregularities, contributing to its extensive inflectional diversity. Recently, advances in Large Language Models (LLMs) have transformed the paradigm in many domains. Despite the success in other fields, current LLMs often fall short in catering to the needs of domain experts like Tibetans, and the potential of LLMs for Tibetan culture is under-explored. The intrinsic reasons are the immense and intricate nature of Tibetan culture as well as the necessity for higher granularity and richness in knowledge. Simultaneously, the complexity and uniqueness of its grammatical structure, coupled with its status as a minority ethnic language, contribute to data scarcity, which remains a fundamental challenge. To alleviate these issues, we introduce Llama-Sunshine (Sun-Shine), the first large language model for Tibetan culture, which is expert in various Tibetan language processing tasks. Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features. We also propose TIB-STC, a comprehensive dataset comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale dataset for Tibetan culture. Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis. Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities. Code and data are avaliable: https://github.com/Vicentvankor/sun-shine.
1 Introduction
Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT Minaee et al. (2024), Gemini Team et al. (2023), Qwen-2.5 Yang et al. (2024a; b), Llama Touvron et al. (2023); Dubey et al. (2024) and DeepSeek Bi et al. (2024); Guo et al. (2025). Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges Achiam et al. (2023). Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA Touvron et al. (2023) and DeepSeek Bi et al. (2024); Guo et al. (2025), which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
Despite their many strengths, LLMs are not tailored to specific fields. Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain Wang et al. (2023a); Zhou et al. (2023). Also, there are some biases against LLMs GLM et al. (2024); Wang et al. (2023b); Blodgett et al. (2020). LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage Achiam et al. (2023). For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments Brown et al. (2020). Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability. Overall, data scarity is the key issue for these areas. Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem Liang et al. (2024). In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language Khysru et al. (2021). Assembling a dataset about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort. Only then can we develop a more specialized and unbiased LLM model on this basis.
To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus dataset designed to address the issue of data scarcity in Tibetan language research. This dataset spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan dataset in the world. On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture. the first LLM specifically tailored for Tibetan language processing. During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 Guo et al. (2025) totally, showing a higher level of knowledge expertise in Tibetan Culture. Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture. Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and dataset for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
In summary, our contributions can be summarized as follows:
  • We present Sun-Shine, the first LLM for Tibetan culture, leveraging Chinese-to-Tibetan translation and manual annotation to tackle linguistic complexities and enhance training accuracy, paving the way for NLP in low-resource languages.
  • We introduce TIB-STC, the first large-scale Tibetan dataset, addressing data scarcity in Tibetan language research. Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
  • Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our dataset, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
2 Related Work
2.1 Tibetan Language Processing
LLMs for rare or low-resource languages are relatively uncommon Gurgurov et al. (2024). The absence of corpus data is the primary and most critical issue Brown et al. (2020); Liang et al. (2024) like Tibetan. However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently Institute (2014). As a result, digital Tibetan language resources are exceptionally scarce. This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing Liang et al. (2024); Khysru et al. (2021).
For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Appendix Figure 7. As shown in Appendix Figure 6, the native Tibetan syllable structure, includes seven kinds of letters. There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel. Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data. To solve this issue, we propose the first large-scale Tibetan dataset, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
2.2 Large Language Model
LLMs, including GPT Achiam et al. (2023), Gemini Team et al. (2023), and ChatGLM GLM et al. (2024); Wang et al. (2023b), are highly parameterized models with billions of parameters, pre-trained on massive natural language datasets Chiang & Lee (2023). Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities. By applying task-specific instructions and fine-tuning techniques Hu et al. ; Rafailov et al. (2024); Yu et al. (2021), LLMs can be optimized for specialized fields, enhancing their performance and domain expertise Zhou et al. (2023).
Taking into account the unique characteristics of Tibetan culture and in order to make life more convenient for the Tibetan people, we introduce the first LLM for Tibetan, Sun-Shine which is based on TIB-STC and cover a wide range of subjects, such as medicine, law, sports, daily life and basic sciences, providing adequate support can significantly improve the daily lives of people in Tibetan areas.
3 TIB-STC and Sun-Shine
3.1 Dataset: TIB-STC
TIB-STC, is the first comprehensive and detailed Tibetan corpus dataset for LLM, consisting of five categories of data as shown in the dataset branch of Table 1. Corpus is used for LLM pre-training, while other datasets focus on optimization and enhancement. For these four sub-datasets, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability. Pre-Training involves the largest dataset in terms of instances, tokens, and storage, as it focuses on general language model training. In the stage of Supervised Fine-Tuning (SFT), datasets are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety). Direct Preference Optimization (DPO) uses the smallest datasets but focuses on high-quality preference data for alignment and optimization.
Dataset Instance Tokens Storage Size
Corpus 122,756,184 11B 18GB
Alpaca-Ti 42,676 76M 109MB
Safety-Prompts-Ti 38,156 62M 116MB
CValues-Ti 2,000 6M 9MB
hh-rlhf-Ti 47,411 98M 154MB
Table 1: The subcategories of TIB-STC and their data size statistics.
For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language. Unlike modern Tibetan, this dataset preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
3.1.1 Corpus
Corpus for Sun-Shine’s pre-training consists of literature (66%), web pages (24%), and media (10%). Content distribution is led by philosophy (30%), medicine (24%), and culture (18%), with smaller portions from academic papers (10%), news (9%), and law (9%). Incorporating book texts enhances linguistic depth and strengthens long-term semantic modeling in LLMs. There are 122,756,184 such instances. Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation. Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base. As shown in Appendix Figure 8, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language. Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
3.1.2 Alpaca-Ti
Alpaca-Ti is used for instruction fine-tuning. We draw inspiration from Alpaca Taori et al. (2023a; b) project at Stanford University and develop the first Tibetan dataset specifically designed for self-supervised instruction fine-tuning in the world. It covers the following topics as shown in Table 2.
Domain Category
Daily Conversation Small Talk and Q&A
Education Multidisciplinary Problem and Solving
Coding Code Generation and Debugging
Creative Writing Poetry, Story
Practical Tasks Summary, Translation, Text Proofreading
Scientific Reasoning Logical Reasoning and Calculation
Areas of Expertise Medicine, Law, Finance, etc.
Entertainment Riddles, Jokes, Game Suggestions.
Table 2: Details of Alpaca-Ti.
Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities. It has a total of 42,676 instances. One data sample is shown in Appendix Figure 9.
3.1.3 Safety-Prompts-Ti
Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table 3, and the data size is 38,156. LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts Chao et al. (2024); Zheng et al. (2024); Sun et al. (2023). Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a dataset to simulate the process in advance. One data sample is shown in Appendix Figure 10.
Instruction Attacks 15k
Instruction Attacks consists of Goal Hijacking 2.5K
Prompt Leaking 2.5K
RolePlay Instruction 2.5K
Unsafe Instruction Topic 2.5K
Inquiry with Unsafe Opinion 2.5K
Reverse Exposure 2.5K
Typical Safety Scenarios 23K
Typical Safety Scenarios consists of Insult 3.6K
Unfairness and Discrimination 3.8K
Crimes and Illegal Activities 3.6K
Physical Harm 1.7K
Mental Health 4K
Privacy and Property 1.3K
Ethics and Morality 1.3K
Others 4.7K
Table 3: Statistical details of Instruction Attacks and Typical Safety Scenarios.
3.1.4 CValues-Ti
CValues-Ti is the dataset, used to improve the security of Tibetan language models. Considering work Moore et al. (2024); Rozen et al. (2024); Xu et al. (2023), we propose this branch of a small dataset because LLM can pose risks or has negative social impacts. For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response). Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe). Unlike the work Xu et al. (2023), We streamlined and condensed the instructions, and compared to their dataset (20K), our sub-dataset, CValues-Ti (2K), achieved equivalent results in subsequent experiments. One data sample is shown in Appendix Figure 11.
3.1.5 hh-rlhf-Ti
hh-rlhf-Ti is also the first dataset that improves the model effect by introducing human intervention. It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF Bai et al. (2022); Kirk et al. , but we reset and summarize the previous dataset based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives. This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan. It has a total of 47,411 instances, and one data sample is shown in Appendix Figure 12.
3.2 Sun-Shine
Figure 1: The training process of the Sun-Shine, consists of three key stages: Pre-training, SFT, and DPO. Pretraining endows the Sun-Shine with a robust foundation in language understanding and generative tasks. SFT refines the Sun-Shine’s ability to interpret specific instructions and generates outputs that adhere to safety standards. DPO adjusts the Sun-Shine to prioritize user preferences, resulting in outputs that are more aligned and appealing.
Sun-Shine is built on the LLaMA 3 Touvron et al. (2023); Dubey et al. (2024) architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community. For more accessible training, we adopt the LLaMA 3.1 (8B) model Dubey et al. (2024). Additionally, we integrate NLP model translation with human annotation Üstün et al. (2024) as supplementary data to enhance Sun-Shine’s accuracy.
As illustrated in Figure 1, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO). For training details, Appendix Table 6 has shown parameters’ setting. Notably, we adopted LoRA-based parameter-efficient fine-tuning Hu et al. for both SFT and DPO, which significantly reduced memory usage without sacrificing performance.
3.2.1 Pre-Training
During the Pre-training stage, the Sun-Shine is pre-trained on Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction. And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text. For details,
3.2.2 Supervised Fine-Tuning
Following pretraining, the model undergoes supervised fine-tuning Yu et al. (2021) with Alpaca-Ti and Safety-Prompts-Ti. During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
3.2.3 Direct Preference Optimization
In the stage of DPO Rafailov et al. (2024); Chao et al. (2024); Zheng et al. (2024), the model’s alignment and output quality are further optimized using preference data. The input data comes from CValues-Ti and hh-rlhf-Ti. Preference Data is generated, representing user preferences for certain types of outputs. The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
4 Evaluation: TLUE Benchmark
Unfortunately, existing benchmarks Li et al. (2024); Zhang et al. (2024); Sun et al. (2023) are available only in Chinese or English, with no purely Tibetan versions. To address this issue, we first translated the data and manually corrected them. This has been made into a new benchmark, TLUE Gao et al. (2025). Fortunately, our partnerships include Tibetan institutions based in Tibet. Unlike modern Tibetan, Our translation conversion strives to preserve the ancient grammar and unique linguistic features of traditional Tibetan culture. All data have been carefully validated by Tibet experts and reviewed by relevant institutions. TLUE has two branches: Ti-MMLU and Ti-Safetybench. In this paper, we use both of them to assess Sun-Shine’s Tibetan language ability.
4.1 Ti-MMLU
CMMLU Li et al. (2024) is a comprehensive Chinese benchmark that spans multiple disciplines, including natural sciences, social sciences, engineering, and the humanities, which consists of 11,528 questions distributed across 67 subjects, with each subject containing a minimum of 105 questions. These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions. The benchmark encompasses 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains. Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere. The evaluation matrix includes Response Rate, Accuracy, and Conditional Accuracy. Response rate measures the proportion of instances where LLMs provide a valid response. Accuracy indicates how often LLMs’ selected answer matches the correct one, while conditional accuracy assesses LLMs’ accuracy only when a valid response is given.
4.2 Ti-Safetybench
SafetyBench Zhang et al. (2024); Sun et al. (2023) comprises 11,435 diverse multiple choice questions spanning across 7 distinct categories of safety concerns: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality and Privacy and Property. Its sources are diverse, comprising questions from existing Chinese and English datasets, safety-related examinations, and ChatGPT-augmented samples. Although our prompts ask for direct answers, many models automatically analyze each option. To calculate these metrics, we use two methods: Direct Answer Calculation (DA) and Concern All Answer Calculation (CAA). DA evaluates the model’s ability to respond to Tibetan instructions by checking if its selected answer is valid and correct. CAA assesses the performance ceiling by filtering valid choices, removing full answer combinations, and verifying correctness based on the remaining choice.
5 Experimental Result
We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) Achiam et al. (2023), Llama-3.18B Dubey et al. (2024); Touvron et al. (2023), Qwen-2.5 (72B) Yang et al. (2024a) and DeepSeek-R1 Bi et al. (2024); Guo et al. (2025). Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis.
5.1 Insights from Performance Results
Ti-MMLU: as shown in Figure 2, Figure 3 and Appendix Table 7, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs .
(a) Response Rate
(b) Conditional Accuracy
(c) Accuracy
Figure 2: Ti-MMLU benchmark comparison: Sun-Shine vs. other LLMs.
(a) The highest total number of indicators.
(b) The second total number of indicators.
Figure 3: Statistics of Sun-Shine ranks first and second based on the number of indicators, across all categories of the Ti-MMLU benchmark.
For Response Rate, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1 (see Appendix Table 8 and Figure 3). Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training dataset.
For Conditional Accuracy, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead (Appendix Table 9). Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities. Sun-Shine excels among most LLMs but lags slightly behind DeepSeek-R1 due to the larger dataset used to train the latter’s base model.
For Accuracy, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas. This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient. Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance. Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine. Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
Ti-Safetybench: as shown in Figure 4 and Appendix Table 11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods. Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA. GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
(a) Comparison of LLMs Across DA
(b) Comparison of LLMs Across CCA
Figure 4: Ti-SafetyBench comparison: Sun-Shine vs. other LLMs.
Overall, Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1. While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance. This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
5.2 Sun-Shine for Tibetan Culture
As shown in Appendix Figure 13, Appendix Figure 14, and Appendix Figure 15, Sun-Shine maintains a neutral tone when addressing common questions, avoiding exaggerated praise or references to Tibetan classics. In contrast, GPT-4o adopts a more amplified and majestic style, reminiscent of divine scripture, while DeepSeek-R1’s responses frequently allude to the bodhisattva Mañjuśrī, appearing slightly ”overfitted” with more refined yet noticeably embellished language than GPT-4o.
Figure 5: Experimental results of classical Chinese’s output from Sun-Shine, GPT-4o and DeepSeek-R1.
As shown in Figure 5, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese. This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records. The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
Overall, Sun-Shine surpasses DeepSeek-R1 (671B) and GPT-4o in most performance indicators, with its classical Tibetan style attributed to the curated TIB-STC and training strategy, where human annotation and correction played a key role.
5.3 Downstream Task
Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan). BELU Papineni et al. (2002) and ROUGE Lin (2004) are used to evaluate this task. As shown in Table 4, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower. Table 5 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
LLM B-4 R-1 R-2 R-l
Sun-Shine 42.16 65.78 48.91 51.58
GPT-4o 38.85 61.12 44.33 47.02
DeepSeek-R1 41.85 65.49 47.73 51.04
Qwen-2.5 35.20 58.59 41.25 44.73
Table 4: Comparison of translation task indicators with other LLMs.
Tibetan Language Proficiency Test: we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
LLM A B C D
Sun-Shine 63 61 44 42
GPT-4o 45 47 53 51
DeepSeek-R1 59 56 67 66
Qwen-2.5 41 43 53 54
Table 5: Comparison of Tibetan Language Proficiency Test with other LLMs.
As shown in Table 5, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics. This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our dataset, TIB-STC, which includes a richer collection of ancient Tibetan texts.
6 Conclusion
In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC dataset, marking significant progress in Tibetan language processing. Sun-Shine, trained on the comprehensive TIB-STC dataset, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios. TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data. Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan. Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our dataset, TIB-STC, in providing a high-quality ancient Tibetan corpus. This work advances low-resource language models and underscores the importance of culturally specific datasets.
Acknowledgments
This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.
Appendix A Appendix A
A.1 Tibetan Components and their Composition
Figure 6: This is the native tibetan syllable structure, which means raise or lift. It consists of 7 parts: prefix letter, root letter, superscript letter, subscript letter, vowel letter, suffix letter, and farther suffix letter.
Figure 7: Characters belong to each component.
A.2 Details of TIB-STC
A.2.1 Corpus Subset
Figure 8: The two data samples from the corpus of TIB-STC are from the categories of ”Buddhism” and ”Culture” respectively.
A.2.2 Fine-Tuning Subset: Alpaca-Ti
Figure 9: One sample from Alpaca-Ti and ’-’ means nothing was entered.
A.2.3 Fine-Tuning Subset: Safety-Prompts-Ti
Figure 10: One sample from Safety-Prompts-Ti.
A.2.4 Alignment Subset: CValues-Ti
Figure 11: One sample from CValues-Ti.
A.2.5 Alignment Subset: hh-rlhf-Ti
Figure 12: One sample from hh-rlhf-Ti and ’…’ means content where a person interacts with an assistant that the assistant represents the LLM. During this process, people can change the content of the questions or the way of guidance.
A.3 Hyperparameter Settings for Each Stage of Sun-Shine Training
Parameter Training Stage
Pre-Training SFT DPO
Precision bfloat16 bfloat16 bfloat16
Optimizer AdamW AdamW AdamW
Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $5\times 10^{-6}$
LR Scheduler CosineAnnealing CosineAnnealing CosineAnnealing
Warmup Ratio 0.1 0.1 0.1
Gradient Accumulation 4 4 8
Epochs 3 3 3
Batch Size (tokens) 4096 4096 4096
Max Seq Length 4096 4096 4096
LoRA No Yes (rank=32) Yes (rank=32)
Weight Decay 0.1 0.1 0.1
Dropout 0.1 0.1 0.1
pref_beta 0.1
pref_loss Sigmoid
Table 6: Training hyperparameters used in each stage of Sun-Shine’s development. Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages. LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO. All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1. The pretraining phase involved over 11B tokens with three full epochs. This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
Appendix B Appendix
B.1 Details of Experimental Results on TLUE Benchmark
Note: the highest-ranking; the second-highest) (BLEU-4 = B-4; ROUGE-1 = R-1; ROUGE-2 = R-2; ROUGE-l = R-l) S-S = Sun-Shine; G-4o = GPT-4o; L-318B = Llama-3.18B; G-35T = GPT-3.5-Turbo; Q-2.5 = Qwen-2.5; Go1-m = GPT-o1-mini; DK-r1 = DeepSeek-r1; Offensiveness = OFF, Unfairness and Bias = UB, Physical Health = PH, Mental Health = MH, Illegal Activities = IA, Ethics and Morality = EM and Privacy and Property = PP.
B.1.1 Overview of LLMs Experiments on the Branch of TLUE: Ti-MMLU
Model Response Rate
China Specific Social Sciences Humanities STEM Other
Sun-Shine 0.5507 0.5588 0.5377 0.5481 0.5498
GPT-4o 0.4932 0.5133 0.5107 0.4915 0.4957
Llama-3.18B 0.1907 0.1998 0.2056 0.2082 0.2104
GPT-3.5-turbo 0.0687 0.0741 0.0748 0.0806 0.0801
Qwen-2.5 0.2563 0.2703 0.2736 0.2441 0.2419
GPT-o1-mini 0.2050 0.2340 0.2333 0.2231 0.2241
DeepSeek-R1 0.4206 0.4069 0.3884 0.3580 0.3482
Model Conditional Accuracy
China Specific Social Sciences Humanities STEM Other
Sun-Shine 0.3208 0.3367 0.3477 0.3457 0.3489
GPT-4o 0.2845 0.2871 0.2846 0.2746 0.2775
Llama-3.18B 0.2510 0.2704 0.2662 0.2620 0.2597
GPT-3.5-turbo 0.2489 0.2605 0.2656 0.2710 0.2731
Qwen-2.5 0.2656 0.2655 0.2710 0.2866 0.2900
GPT-o1-mini 0.2636 0.2700 0.2716 0.2742 0.2728
DeepSeek-R1 0.3209 0.3482 0.3524 0.3584 0.3573
Model Accuracy
China Specific Social Sciences Humanities STEM Other
Sun-Shine 0.1888 0.1742 0.1683 0.1614 0.1618
GPT-4o 0.1594 0.1676 0.1634 0.1515 0.1535
Llama-3.18B 0.0656 0.0628 0.0723 0.0662 0.0663
GPT-3.5-turbo 0.0182 0.0201 0.0205 0.0277 0.0265
Qwen-2.5 0.0691 0.0723 0.0744 0.0702 0.0709
GPT-o1-mini 0.0475 0.0584 0.0598 0.0603 0.0606
DeepSeek-R1 0.1863 0.1819 0.1705 0.1562 0.1534
Table 7: Benchmarking Sun-Shine Against Leading LLMs on Ti-MMLU . For more detailed information about these three indicators, please refer to Appendix Table 8 (Response Rate), Appendix Table 9 (Conditional Accuracy) and Appendix Table 10 (Accuracy). The Category column lists the categories in order from top to bottom: China Specific, Social Sciences, Humanities, STEM, and Other.
B.1.2 Details of Response Rate
Category Response Rate
S-S G-4o L-318 G-35T Q-2.5 Go1-m DK-r1
ancient_chinese 0.5366 0.4695 0.1402 0.0732 0.2500 0.2317 0.3232
chinese_civil_service_exam 0.4938 0.4062 0.1688 0.0438 0.2312 0.1000 0.3812
chinese_food_culture 0.5588 0.5147 0.1250 0.0809 0.2279 0.1765 0.4191
chinese_foreign_policy 0.5206 0.4981 0.3458 0.0093 0.2804 0.1589 0.4579
chinese_teacher_qualification 0.5972 0.6089 0.2682 0.0726 0.2458 0.3073 0.4078
elementary_chinese 0.5238 0.4167 0.1667 0.0556 0.2817 0.1667 0.4683
ethnology 0.5481 0.5037 0.1852 0.0963 0.3259 0.2148 0.6148
high_school_politics 0.5105 0.4955 0.1888 0.0979 0.2098 0.1748 0.4685
modern_chinese 0.5431 0.5241 0.1724 0.0862 0.2069 0.2328 0.3707
chinese_literature 0.6127 0.5412 0.1618 0.0833 0.2892 0.2500 0.5000
chinese_history 0.5635 0.4489 0.1672 0.0526 0.2632 0.1734 0.3839
chinese_driving_rule 0.5198 0.4420 0.1985 0.0687 0.2519 0.2595 0.3817
construction_project_management 0.6619 0.4388 0.1942 0.0647 0.2662 0.2230 0.3453
elementary_commonsense 0.5404 0.5354 0.1414 0.0758 0.2828 0.1465 0.4848
traditional_chinese_medicine 0.5459 0.5405 0.2378 0.0703 0.2324 0.2595 0.3027
business_ethics 0.5359 0.5167 0.2057 0.0574 0.2919 0.2249 0.3780
economics 0.5283 0.4465 0.2767 0.0440 0.2516 0.1824 0.3396
education 0.5460 0.5215 0.1840 0.0920 0.2761 0.3129 0.4049
college_education 0.5981 0.5234 0.2336 0.1589 0.4299 0.2617 0.4673
journalism 0.5947 0.5447 0.1744 0.0407 0.3256 0.2209 0.4826
marketing 0.6556 0.5944 0.2278 0.0889 0.2667 0.3500 0.2389
professional_accounting 0.6114 0.5829 0.1829 0.0800 0.1943 0.2400 0.3714
professional_psychology 0.5645 0.5303 0.2198 0.0819 0.1940 0.2586 0.3578
public_relations 0.5920 0.5172 0.1437 0.0862 0.2299 0.2931 0.4598
security_study 0.5704 0.4963 0.1630 0.1185 0.4074 0.3037 0.4000
high_school_geography 0.5678 0.4831 0.2712 0.0678 0.3305 0.1780 0.4237
management 0.5905 0.5857 0.2524 0.0857 0.2429 0.3190 0.3810
sociology 0.5442 0.5177 0.1991 0.0442 0.2832 0.3319 0.3805
marxist_theory 0.6085 0.5556 0.2804 0.0529 0.3122 0.2804 0.4497
college_law 0.4648 0.4519 0.2315 0.0833 0.2870 0.2500 0.3981
global_facts 0.5235 0.4930 0.2282 0.0738 0.3423 0.2349 0.4027
international_law 0.5568 0.5081 0.2270 0.0595 0.3351 0.2541 0.0919
jurisprudence 0.5377 0.5109 0.3187 0.0681 0.3358 0.2409 0.3869
world_religions 0.5000 0.4938 0.1125 0.0938 0.2625 0.1812 0.4688
logical 0.5366 0.4715 0.1463 0.0976 0.2195 0.2195 0.2846
professional_law 0.5071 0.4882 0.2654 0.0474 0.2180 0.2180 0.3033
philosophy 0.5524 0.4667 0.2476 0.0952 0.2190 0.2286 0.0476
world_history 0.5342 0.4907 0.2174 0.0683 0.3354 0.2298 0.4720
arts 0.0750 0.5437 0.1500 0.1000 0.2375 0.2125 0.4500
electrical_engineering 0.5698 0.5349 0.2035 0.1337 0.2965 0.2384 0.3663
actuarial_science 0.3692 0.1887 0.1698 0.1792 0.0642 0.1038 0.5000
engineering_hydrology 0.5755 0.5377 0.2170 0.0660 0.3113 0.3208 0.5189
genetics 0.5455 0.4830 0.1932 0.0795 0.1648 0.1761 0.2216
astronomy 0.5758 0.4606 0.2424 0.0727 0.1455 0.2182 0.3879
college_mathematics 0.5048 0.3238 0.2095 0.1429 0.1619 0.1524 0.1429
college_medical_statistics 0.6038 0.5094 0.1698 0.0566 0.1792 0.1887 0.0660
virology 0.5680 0.5089 0.1479 0.0414 0.1893 0.2899 0.2071
computer_science 0.5686 0.4216 0.1961 0.0882 0.1225 0.1814 0.2696
conceptual_physics 0.5986 0.5918 0.3878 0.0340 0.1565 0.2925 0.2109
anatomy 0.6622 0.5203 0.2027 0.0676 0.1689 0.2635 0.2905
machine_learning 0.5656 0.4754 0.1967 0.0574 0.1639 0.1393 0.2951
high_school_physics 0.6364 0.5455 0.1909 0.0727 0.2455 0.2198 0.3415
high_school_mathematics 0.5915 0.2317 0.1646 0.2012 0.1341 0.1098 0.3415
high_school_chemistry 0.5758 0.3409 0.2955 0.0682 0.1591 0.1288 0.0985
high_school_biology 0.6746 0.4852 0.2604 0.0828 0.1657 0.2249 0.0947
elementary_mathematics 0.5996 0.3870 0.1913 0.1565 0.1696 0.1478 0.5478
legal_and_moral_basis 0.5748 0.5561 0.2523 0.0888 0.3131 0.2523 0.5514
computer_security 0.5789 0.5088 0.2398 0.1053 0.2105 0.2339 0.2164
food_science 0.4895 0.4705 0.1469 0.0909 0.2587 0.1818 0.3077
college_medicine 0.4982 0.4908 0.2674 0.0366 0.1722 0.1538 0.1758
clinical_knowledge 0.5738 0.4557 0.1899 0.1013 0.2447 0.2616 0.1308
professional_medicine 0.5798 0.5505 0.3165 0.0771 0.2527 0.2287 0.2074
human_sexuality 0.5635 0.5317 0.2302 0.0317 0.2778 0.2143 0.3651
agronomy 0.5621 0.5089 0.1775 0.0651 0.2071 0.2485 0.4201
sports_science 0.6424 0.5697 0.2000 0.0909 0.2848 0.2667 0.3879
nutrition 0.5586 0.4759 0.1862 0.0621 0.1172 0.2414 0.2552
elementary_IT 0.5462 0.5378 0.2353 0.1050 0.1975 0.2353 0.2647
Table 8: Response Rate Comparison: Sun-Shine vs. Other LLMs.
B.1.3 Details of Conditional Accuracy
Category Conditional Accuracy
S-S G-4o L-318 G-35T Q-2.5 Go1-m DK-r1
ancient_chinese 0.2697 0.2500 0.2174 0.2500 0.2195 0.2812 0.2500
chinese_civil_service_exam 0.2562 0.2152 0.2222 0.1429 0.3243 0.2368 0.2658
chinese_food_culture 0.3258 0.3143 0.3529 0.2727 0.3226 0.2143 0.3171
chinese_foreign_policy 0.3444 0.3438 0.2973 0.0000 0.2667 0.3235 0.3448
chinese_teacher_qualification 0.2921 0.2752 0.1667 0.3077 0.3636 0.2571 0.2650
elementary_chinese 0.3126 0.2813 0.2857 0.4286 0.2958 0.2533 0.3145
ethnology 0.3719 0.4109 0.2400 0.2308 0.3864 0.3333 0.5714
high_school_politics 0.3662 0.3225 0.2963 0.2143 0.1667 0.2174 0.3412
modern_chinese 0.3592 0.3265 0.2000 0.4000 0.1250 0.1852 0.3489
chinese_literature 0.2980 0.2000 0.3030 0.1765 0.3220 0.3077 0.2879
chinese_history 0.3407 0.2552 0.2778 0.2941 0.2353 0.2418 0.2543
chinese_driving_rule 0.3299 0.2764 0.1923 0.1111 0.2424 0.3247 0.3721
construction_project_management 0.3470 0.2951 0.2593 0.2222 0.2162 0.2292 0.3253
elementary_commonsense 0.3017 0.2925 0.2500 0.4667 0.1964 0.2759 0.2654
traditional_chinese_medicine 0.2971 0.2100 0.2045 0.2308 0.3023 0.2738 0.2912
business_ethics 0.3981 0.3411 0.2791 0.3333 0.2623 0.2391 0.3802
economics 0.3597 0.3476 0.2955 0.2857 0.1500 0.3341 0.3254
education 0.3667 0.3421 0.3647 0.2667 0.3111 0.1818 0.3845
college_education 0.3125 0.2143 0.3600 0.2353 0.2174 0.2857 0.3478
journalism 0.2981 0.2366 0.2667 0.2857 0.2679 0.2281 0.3098
marketing 0.3768 0.3051 0.2683 0.2500 0.3750 0.3541 0.3254
professional_accounting 0.3748 0.3084 0.3438 0.5714 0.2647 0.3451 0.3746
professional_psychology 0.3985 0.2823 0.2157 0.3158 0.3333 0.3698 0.3945
public_relations 0.2427 0.1889 0.3600 0.1333 0.2250 0.2845 0.4752
security_study 0.2997 0.3333 0.3182 0.3125 0.3636 0.2258 0.3977
high_school_geography 0.4179 0̆.2807 0.2188 0.1250 0.2308 0.2368 0.3627
management 0.3089 0.2913 0.2264 0.3333 0.2941 0.3154 0.5369
sociology 0.3333 0.2997 0.2889 0.1000 0.1562 0.2062 0.3221
marxist_theory 0.3217 0.3143 0.2453 0.1000 0.2712 0.3210 0.3212
college_law 0.3287 0.2842 0.2800 0.2222 0.2258 0.2165 0.3109
global_facts 0.3636 0.3599 0.3235 0.0909 0.2353 0.3000 0.3611
international_law 0.3830 0.2524 0.2381 0.3636 0.2903 0.2714 0.3799
jurisprudence 0.3258 0.2667 0.2519 0.2857 0.2754 0.2786 0.3299
world_religions 0.5975 0.2911 0.2222 0.2667 0.2857 0.2143 0.6475
logical 0.2931 0.2879 0.2778 0.2500 0.2593 0.2849 0.3648
professional_law 0.3178 0.2223 0.1786 0.2000 0.2609 0.2865 0.3802
philosophy 0.3276 0.2429 0.2308 0.3000 0.3478 0.3169 0.3057
world_history 0.5658 0.2907 0.2286 0.5455 0.3148 0.3267 0.4407
arts 0.2759 0.2500 0.3333 0.4375 0.3684 0.2157 0.2453
electrical_engineering 0.3349 0.3478 0.3114 0.1304 0.2745 0.2456 0.2857
engineering_hydrology 0.3684 0.1148 0.1304 0.5714 0.2121 0.2500 0.4730
astronomy 0.3903 0.2237 0.2750 0.3750 0.3333 0.2111 0.3699
medical_statistics 0.2978 0.2656 0.2222 0.5000 0.2105 0.3636 0.3026
computer_science 0.4013 0.2845 0.2500 0.2778 0.3600 0.3966 0.3923
conceptual_physics 0.2729 0.2473 0.2712 0.2000 0.2174 0.2632 0.2799
anatomy 0.2410 0.2338 0.1333 0.2000 0.2000 0.1667 0.2367
machine_learning 0.2319 0.2241 0.2083 0.1429 0.2000 0.2498 0.2729
high_school_physics 0.2667 0.1857 0.1429 0.1250 0.2593 0.2441 0.6366
high_school_mathematics 0.6667 0.5000 0.3333 0.4500 0.5268 0.5246 0.5241
high_school_chemistry 0.2667 0.1974 0.3590 0.1111 0.2857 0.2459 0.3636
high_school_biology 0.3171 0.2544 0.2727 0.5000 0.2500 0.2321 0.3295
elementary_mathematics 0.3146 0.2519 0.2273 0.2778 0.4615 0.2830 0.2941
actuarial_science 0.3400 0.2500 0.4444 0.1579 0.3214 0.2800 0.4737
genetics 0.2941 0.2555 0.3529 0.2143 0.3793 0.3011 0.3582
college_mathematics 0.5882 0.2642 0.2727 0.2000 0.4941 0.1905 0.4896
virology 0.2083 0.1744 0.1200 0.4286 0.2188 0.3094 0.2157
legal_and_moral_basis 0.3821 0.3277 0.2963 0.1579 0.3284 0.2436 0.3684
computer_security 0.3223 0.2848 0.3171 0.1667 0.3167 0.2750 0.3033
food_science 0.3914 0.3824 0.1905 0.2308 0.2703 0.2250 0.4154
college_medicine 0.3134 0.2500 0.2055 0.2000 0.2128 0.2022 0.3326
clinical_knowledge 0.3448 0.3088 0.2222 0.3333 0.3426 0.2143 0.3248
professional_medicine 0.2431 0.2222 0.2605 0.3448 0.2737 0.2569 0.3433
human_sexuality 0.3134 0.3099 0.1726 0.7500 0.3714 0.3174 0.2989
agronomy 0.2791 0.2558 0.2667 0.0909 0.3714 0.2188 0.2626
sports_science 0.4171 0.2666 0.3936 0.2667 0.4465 0.2969 0.3830
nutrition 0.2319 0.2222 0.1481 0.2222 0.2941 0.3833 0.4412
elementary_IT 0.3185 0.3906 0.2500 0.3600 0.2128 0.2874 0.3452
Table 9: Conditional Accuracy Comparison: Sun-Shine vs. Other LLMs.
B.1.4 Details of Accuracy
Category Accuracy
S-S G-4o L-318 G-35T Q-2.5 Go1-m DK-r1
ancient_chinese 0.1341 0.1220 0.3050 0.0183 0.0549 0.0549 0.1159
chinese_civil_service_exam 0.1062 0.1000 0.0375 0.0062 0.0750 0.0025 0.1562
chinese_food_culture 0.1765 0.1618 0.0441 0.0221 0.0735 0.0294 0.1838
chinese_foreign_policy 0.2056 0.2028 0.1028 0.0000 0.0748 0.0654 0.1869
chinese_teacher_qualification 0.1676 0.1553 0.0447 0.0223 0.0894 0.0503 0.2458
elementary_chinese 0.1866 0.1695 0.0476 0.0238 0.0833 0.0397 0.1865
ethnology 0.2296 0.2456 0.0444 0.0222 0.1259 0.0519 0.3333
high_school_politics 0.1738 0.1834 0.0559 0.0210 0.0350 0.0210 0.2098
modern_chinese 0.1897 0.1654 0.0345 0.0345 0.0259 0.0603 0.1466
chinese_literature 0.1765 0.1882 0.0490 0.0147 0.0931 0.0686 0.1961
chinese_history 0.1920 0.1146 0.0464 0.0155 0.0619 0.0372 0.1238
chinese_driving_rule 0.1676 0.1492 0.0382 0.0076 0.0611 0.0763 0.1985
construction_project_management 0.2230 0.1295 0.0504 0.0144 0.0576 0.0504 0.1727
elementary_commonsense 0.1966 0.1714 0.0354 0.0354 0.0556 0.0354 0.2424
traditional_chinese_medicine 0.1568 0.1323 0.0486 0.0162 0.0703 0.0703 0.0973
business_ethics 0.2057 0.1852 0.0574 0.0191 0.0766 0.0478 0.1914
economics 0.1195 0.1195 0.0818 0.0126 0.0377 0.0252 0.1824
education 0.1865 0.1795 0.0675 0.0245 0.0859 0.0613 0.1779
college_education 0.1867 0.2100 0.0741 0.0374 0.0935 0.1028 0.2056
journalism 0.1829 0.1676 0.0465 0.0116 0.0872 0.0640 0.2035
marketing 0.1833 0.2000 0.0611 0.0222 0.1000 0.1167 0.1667
professional_accounting 0.1886 0.1714 0.0629 0.0457 0.0514 0.0629 0.1523
professional_psychology 0.1897 0.1736 0.0474 0.0259 0.0647 0.0690 0.1681
public_relations 0.1437 0.0977 0.0517 0.0115 0.0517 0.0862 0.1389
security_study 0.1333 0.2222 0.0519 0.0370 0.1481 0.0593 0.1259
high_school_geography 0.1356 0.2333 0.0593 0.0085 0.0763 0.0508 0.2034
management 0.1810 0.1695 0.0571 0.0286 0.0714 0.1163 0.2143
sociology 0.1593 0.1726 0.0575 0.0044 0.0442 0.0619 0.1681
marxist_theory 0.1957 0.1799 0.0688 0.0053 0.0847 0.0741 0.2222
college_law 0.1574 0.1222 0.0648 0.0185 0.0648 0.0833 0.1759
global_facts 0.1822 0.1621 0.0738 0.0067 0.0805 0.0671 0.1409
international_law 0.1405 0.1965 0.0541 0.0216 0.0973 0.0595 0.0270
jurisprudence 0.1405 0.1998 0.0803 0.0195 0.0925 0.0706 0.1752
world_religions 0.1437 0.1188 0.0250 0.0250 0.0750 0.0312 0.2750
logical 0.1546 0.1382 0.0407 0.0244 0.0569 0.0813 0.1220
professional_law 0.1611 0.1009 0.0474 0.0095 0.0569 0.0474 0.0758
philosophy 0.0667 0.1810 0.0571 0.0286 0.0762 0.0571 0.0095
world_history 0.1543 0.1304 0.0497 0.0373 0.1056 0.0807 0.1577
arts 0.1880 0.1500 0.5000 0.0438 0.0875 0.0438 0.1750
electrical_engineering 0.1429 0.1190 0.0581 0.0174 0.0814 0.0523 0.2209
engineering_hydrology 0.1981 0.1799 0.0283 0.3770 0.0660 0.0755 0.1792
astronomy 0.1273 0.1030 0.0667 0.0242 0.0545 0.0970 0.1697
medical_statistics 0.1604 0.1415 0.0377 0.0283 0.0377 0.0472 0.0283
computer_science 0.1618 0.1275 0.0049 0.0245 0.0441 0.0441 0.1127
conceptual_physics 0.1497 0.1361 0.1020 0.0068 0.0340 0.0748 0.1224
anatomy 0.2230 0.1216 0.0270 0.0135 0.0338 0.0405 0.0878
machine_learning 0.1311 0.1000 0.0410 0.0082 0.0328 0.0328 0.1066
high_school_physics 0.1818 0.1964 0.0273 0.0091 0.0636 0.0455 0.2091
high_school_mathematics 0.1182 0.1453 0.0549 0.0549 0.1341 0.0671 0.2378
high_school_chemistry 0.1136 0.0970 0.1061 0.0076 0.0455 0.0682 0.0303
high_school_biology 0.1716 0.1534 0.0710 0.0414 0.0414 0.0651 0.0178
elementary_mathematics 0.1432 0.1217 0.0435 0.0435 0.0783 0.0478 0.1139
actuarial_science 0.0755 0.0472 0.0755 0.0283 0.0849 0.0660 0.2453
genetics 0.1307 0.1403 0.0682 0.0170 0.0625 0.0739 0.0625
college_mathematics 0.1333 0.0952 0.0571 0.0286 0.0952 0.0286 0.0762
virology 0.1183 0.0888 0.0178 0.0178 0.0414 0.1183 0.0769
legal_and_moral_basis 0.2182 0.2022 0.0748 0.0140 0.1028 0.0654 0.3411
computer_security 0.2047 0.1579 0.0076 0.0175 0.0877 0.0877 0.1228
food_science 0.1366 0.1936 0.0280 0.0210 0.0699 0.0140 0.1329
college_medicine 0.1538 0.1245 0.0549 0.0073 0.0366 0.0366 0.0623
clinical_knowledge 0.1772 0.1565 0.0422 0.0338 0.0844 0.0591 0.0549
professional_medicine 0.1410 0.1230 0.0824 0.0266 0.0691 0.0718 0.0691
human_sexuality 0.1746 0.1667 0.0397 0.0238 0.1032 0.0635 0.1667
agronomy 0.1420 0.1302 0.0473 0.0059 0.0769 0.0533 0.2071
sports_science 0.2242 0.2189 0.0727 0.0242 0.1091 0.0788 0.1333
nutrition 0.1241 0.1103 0.0276 0.0138 0.0345 0.0759 0.1034
elementary_IT 0.1092 0.2102 0.0588 0.0378 0.0420 0.0798 0.1387
Table 10: Accuracy Comparison: Sun-Shine vs. Other LLMs.
B.2 Overview of LLMs Experiments on the Branch of TLUE: Ti-Safetybench
Model Category
Filter Invaild EM IA MH OFF PH PP UB
Sun-Shine DA 59.20% 0.292 0.379 0.204 0.255 0.258 0.307 0.372
GPT-4o 38.59% 0.144 0.357 0.241 0.199 0.246 0.296 0.358
Llama-3.18B 74.75% 0.105 0.088 0.101 0.097 0.094 0.099 0.109
GPT-3.5-turbo 73.42% 0.085 0.073 0.087 0.086 0.097 0.106 0.123
Qwen-2.5 49.50% 0.199 0.118 0.164 0.198 0.150 0.159 0.371
GPT-o1-mini 71.20% 0.109 0.070 0.095 0.111 0.097 0.095 0.170
DeepSeek-R1 51.14% 0.287 0.373 0.232 0.250 0.297 0.309 0.371
Sun-Shine CCA 16.60% 0.483 0.455 0.507 0.448 0.509 0.336 0.466
GPT-4o 34.55% 0.478 0.429 0.429 0.287 0.255 0.387 0.401
Llama-3.18B 69.72% 0.130 0.101 0.123 0.124 0.119 0.115 0.126
GPT-3.5-turbo 67.75% 0.108 0.082 0.103 0.125 0.114 0.118 0.160
Qwen-2.5 16.61% 0.448 0.413 0.427 0.289 0.259 0.385 0.403
GPT-o1-mini 61.69% 0.254 0.099 0.134 0.163 0.135 0.117 0.231
DeepSeek-R1 28.14% 0.558 0.457 0.500 0.429 0.511 0.339 0.457
Table 11: Benchmarking Sun-Shine Against Leading LLMs on Ti-Safetybench.
B.3 Output of Experimental Results from LLMs
Figure 13: Example of Sun-Shine’s output.
Figure 14: Example of GPT-4o’s output.
Figure 15: Example of DeepSeek-R1’s output.

### v3

@@
-Sun-Shine: A Large Language Model for Tibetan Culture
+Sun-Shine: A Foundation Large Language Model for Tibetan Culture and Heritage
 Abstract
@@
 Sun-Shine incorporates state-of-the-art model architectures optimized for Tibetan’s linguistic features.
-We also propose TIB-STC, a comprehensive dataset comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale dataset for Tibetan culture.
+We also propose TIB-STC, a comprehensive benchmark comprising diverse Tibetan texts such as literature, religious scripts, news, and conversational data, which is also the first large-scale benchmark for Tibetan culture.
 Though comprehensive experiments, Sun-Shine not only demonstrates a higher level of knowledge expertise for Tibetan culture but also gains preliminary embodied intelligence capabilities in Tibetan language processing tasks, like language modeling, text classification, machine translation, and syntactic analysis.
 Moreover, it excels in low-resource scenarios, showcasing strong generalization capabilities.
-Code and data are avaliable: https://github.com/Vicentvankor/sun-shine.
+Index Terms:
+Large Language Model, Foundation Model, Tibetan Language Benchmark, Tibetan Language Processing
 1 Introduction
-Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT Minaee et al.
-(2024), Gemini Team et al.
-(2023), Qwen-2.5 Yang et al.
-(2024a; b), Llama Touvron et al.
-(2023); Dubey et al.
-(2024) and DeepSeek Bi et al.
-(2024); Guo et al.
-(2025).
-Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges Achiam et al.
-(2023).
-Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA Touvron et al.
-(2023) and DeepSeek Bi et al.
-(2024); Guo et al.
-(2025), which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
+Large language models (LLMs) have gained significant attention for their remarkable performance across a wide range of natural language tasks, particularly following the release of ChatGPT [1], Gemini [12], Qwen-2.5 [34, 33], Llama [2, 20] and DeepSeek [37, 38].
+Compared to smaller models, LLMs demonstrate superior generalization across a wide range of natural language processing (NLP) tasks and exhibit unique emergent abilities to solve previously unseen or complex challenges [4].
+Although ChatGPT remains non-open source, open-source communities have actively developed several alternative models with relatively affordable training costs, like LLaMA [2] and DeepSeek [37, 38], which establishes LLMs as promising solutions for real-world applications that demand effective communication and reasoning capabilities.
 Despite their many strengths, LLMs are not tailored to specific fields.
-Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain Wang et al.
-(2023a); Zhou et al.
-(2023).
-Also, there are some biases against LLMs GLM et al.
-(2024); Wang et al.
-(2023b); Blodgett et al.
-(2020).
-LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage Achiam et al.
-(2023).
-For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments Brown et al.
-(2020).
+Pre-training is the strategy that make LLMs become ”experts” in specific domains where LLMs need accurate and domain-specific expert knowledge, such as medical domain [3, 5].
+Also, there are some biases against LLMs [6, 7, 8].
+LLMs often reflect biases in training data, including gender, racial, and cultural biases, regional imbalances favoring English and Western cultures, and uneven domain coverage [4].
+For instance, GPT-3’s generated content was found to contain racial stereotypes in multiple experiments [9].
 Mitigation requires diverse data, bias detection tools, and model adjustments to ensure fairness and reliability.
 Overall, data scarity is the key issue for these areas.
-Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem Liang et al.
-(2024).
-In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language Khysru et al.
-(2021).
-Assembling a dataset about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort.
+Especially for some niche languages, such as Tibetan, data scarcity is the most direct and fundamental problem [11].
+In addition, Tibetan itself has a highly complex structure, which is further compounded by the intricate syntax and grammar of the language [10].
+Assembling a benchmark about such complex but elegant words is already a monumental task, but creating one that is fair and unbiased requires even greater effort.
 Only then can we develop a more specialized and unbiased LLM model on this basis.
-To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus dataset designed to address the issue of data scarcity in Tibetan language research.
-This dataset spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan dataset in the world.
+To address this issue, we introduce the TIB-STC, the first comprehensive Tibetan corpus benchmark designed to address the issue of data scarcity in Tibetan language research.
+This benchmark spans diverse domains, including law, public welfare, Tibetan literature, and Tibetan medicine, making it the first large-scale, detailed, and comprehensive Tibetan benchmark in the world.
 On this basis, we proposed Llama-Sunshine (Sun-Shine), the large language model of Tibetan culture.
@@
 During its development, we incorporated Chinese translations and manual annotation techniques to improve training accuracy and enhance its performance.
-Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 Guo et al.
-(2025) totally, showing a higher level of knowledge expertise in Tibetan Culture.
+Through extensive experiments, Sun-Shine demonstrates superiority for Tibetan culture, even slightly better tham DeepSeek-R1 [38] totally, showing a higher level of knowledge expertise in Tibetan Culture.
 Note that TIB-STC is based on criteria manually evaluated by Tibetan experts, and can accurately reflect the capabilities that LLMs possess in the field of Tibetan culture.
-Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and dataset for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
+Together, Sun-Shine and TIB-STC represent not only pioneering efforts as the first LLM and benchmark for Tibetan but also significant advancements in addressing the broader challenges of building LLMs for minority languages and mitigating bias in underrepresented linguistic contexts.
 In summary, our contributions can be summarized as follows:
   • We present Sun-Shine, the first LLM for Tibetan culture, leveraging Chinese-to-Tibetan translation and manual annotation to tackle linguistic complexities and enhance training accuracy, paving the way for NLP in low-resource languages.
-  • We introduce TIB-STC, the first large-scale Tibetan dataset, addressing data scarcity in Tibetan language research.
+  • We introduce TIB-STC, the first large-scale Tibetan benchmark, addressing data scarcity in Tibetan language research.
 Covering domains like law, welfare, literature, and medicine, it serves as a vital resource for advancing Tibetan NLP.
-  • Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our dataset, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
+  • Experimental results demonstrate that Sun-Shine, trained and fine-tuned on our benchmark, TIB-STC, significantly outperforms other LLMs in Tibetan language processing totally, demonstrating superiority for Tibetan culture.
 2 Related Work
 2.1 Tibetan Language Processing
-LLMs for rare or low-resource languages are relatively uncommon Gurgurov et al.
-(2024).
-The absence of corpus data is the primary and most critical issue Brown et al.
-(2020); Liang et al.
-(2024) like Tibetan.
-However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently Institute (2014).
+LLMs for rare or low-resource languages are relatively uncommon [13].
+The absence of corpus data is the primary and most critical issue [9, 11] like Tibetan.
+However, Tibetan-speaking population in Tibet represents a minority within China, with an even smaller subset capable of reading and writing the language proficiently [16].
 As a result, digital Tibetan language resources are exceptionally scarce.
-This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing Liang et al.
-(2024); Khysru et al.
-(2021).
-For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Appendix Figure 7.
-As shown in Appendix Figure 6, the native Tibetan syllable structure, includes seven kinds of letters.
+This scarcity extends to linguistic corpora, which are critical for the development and training of artificial intelligence models, thereby hindering advancements in AI, like LLMs, for Tibetan language processing [11, 10, 26, 19].
+For Tibetan itself, its syllables are two-dimensionally combined by Tibetan characters and we have sorted out the composition distribution of Tibetan characters, shown in Fig. 1.
+Figure 1: Characters Belong to Each Component
+As shown in Fig.
+2, the native Tibetan syllable structure, includes seven kinds of letters.
 There is a longitudinal structure in syllables which is called the upright unit, consisting of root letter, superscript letter, subscript letter, and vowel.
 Such complexity of Tibetan grammar and structure, combined with its status as a minority language, results in a severe scarcity of Tibetan data.
-To solve this issue, we propose the first large-scale Tibetan dataset, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
+To solve this issue, we propose the first large-scale Tibetan benchmark, TIB-STC, the largest and the most comprehensive one in the world now, advancing Tibetan NLP and offering a foundation for future research and potential applications while serving as a model for developing NLP resources for other minority languages.
+Figure 2: One Example of the Native Tibetan Syllable Structure
 2.2 Large Language Model
-LLMs, including GPT Achiam et al.
-(2023), Gemini Team et al.
-(2023), and ChatGLM GLM et al.
-(2024); Wang et al.
-(2023b), are highly parameterized models with billions of parameters, pre-trained on massive natural language datasets Chiang & Lee (2023).
+LLMs, including GPT [4], Gemini [12], and ChatGLM [6, 7], are highly parameterized models with billions of parameters, pre-trained on massive natural language benchmarks [14].
 Their ability to excel in unseen tasks, given only task-specific instructions, underscores their versatility and advanced capabilities.
-By applying task-specific instructions and fine-tuning techniques Hu et al.
-; Rafailov et al.
-(2024); Yu et al.
-(2021), LLMs can be optimized for specialized fields, enhancing their performance and domain expertise Zhou et al.
-(2023).
+By applying task-specific instructions and fine-tuning techniques [15, 17, 18], LLMs can be optimized for specialized fields, enhancing their performance and domain expertise [5].
 Taking into account the unique characteristics of Tibetan culture and in order to make life more convenient for the Tibetan people, we introduce the first LLM for Tibetan, Sun-Shine which is based on TIB-STC and cover a wide range of subjects, such as medicine, law, sports, daily life and basic sciences, providing adequate support can significantly improve the daily lives of people in Tibetan areas.
 3 TIB-STC and Sun-Shine
-3.1 Dataset: TIB-STC
-TIB-STC, is the first comprehensive and detailed Tibetan corpus dataset for LLM, consisting of five categories of data as shown in the dataset branch of Table 1.
-Corpus is used for LLM pre-training, while other datasets focus on optimization and enhancement.
-For these four sub-datasets, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability.
-Pre-Training involves the largest dataset in terms of instances, tokens, and storage, as it focuses on general language model training.
-In the stage of Supervised Fine-Tuning (SFT), datasets are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety).
-Direct Preference Optimization (DPO) uses the smallest datasets but focuses on high-quality preference data for alignment and optimization.
-Dataset Instance Tokens Storage Size
+3.1 Benchmark: TIB-STC
+TIB-STC, is the first comprehensive and detailed Tibetan corpus benchmark for LLM, consisting of five categories of data as shown in the benchmark branch of Table I.
+TABLE I: Subcategories of TIB-STC and Data Size Statistics
+Benchmark Instance Tokens Storage Size
 Corpus 122,756,184 11B 18GB
@@
 hh-rlhf-Ti 47,411 98M 154MB
-Table 1: The subcategories of TIB-STC and their data size statistics.
+Corpus is used for LLM pre-training, while other benchmarks focus on optimization and enhancement.
+For these four sub-benchmarks, Alpaca-Ti and Safety-Prompts-Ti are used for instruction fine-tuning, and CValues-Ti and hh-rlhf-Ti can improves LLM’s alignment ability.
+Pre-Training involves the largest benchmark in terms of instances, tokens, and storage, as it focuses on general language model training.
+In the stage of Supervised Fine-Tuning (SFT), benchmarks are smaller but still sizable, with a focus on task-specific fine-tuning (e.g., instructions and safety).
+Direct Preference Optimization (DPO) uses the smallest benchmarks but focuses on high-quality preference data for alignment and optimization.
 For TIB-STC, all data selection and cleaning are conducted by experts in the Tibetan language.
-Unlike modern Tibetan, this dataset preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
+Unlike modern Tibetan, this benchmark preserves traditional grammar and syntax as much as possible, minimizing the influence of contemporary culture while maintaining its distinct linguistic and cultural characteristics.
 3.1.1 Corpus
@@
 There are 122,756,184 such instances.
+Figure 3: The two data samples from the corpus of TIB-STC are from the categories of ”Buddhism” and ”Culture” respectively.
 Furthermore, leveraging large-scale web text data during pre-training facilitates the acquisition of diverse linguistic patterns and nuances, thereby improving the models’ proficiency in both natural language comprehension and generation.
 Finally, the inclusion of media data offers the unique advantage of establishing an expansive and reliable media knowledge base.
-As shown in Appendix Figure 8, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language.
+As shown in Fig.
+3, we present a news-based training corpus where Chinese and Tibetan serve as the source languages, and English is provided as the translated language.
 Some of the original Chinese versions and Tibetan texts retain classical Chinese sentence structures, which are different from modern Chinese grammar.
@@
 Alpaca-Ti is used for instruction fine-tuning.
-We draw inspiration from Alpaca Taori et al.
-(2023a; b) project at Stanford University and develop the first Tibetan dataset specifically designed for self-supervised instruction fine-tuning in the world.
-It covers the following topics as shown in Table 2.
+We draw inspiration from Alpaca [31, 32] project at Stanford University and develop the first Tibetan benchmark specifically designed for self-supervised instruction fine-tuning in the world.
+It covers the following topics as shown in Table II.
+TABLE II: Details of Alpaca-Ti
 Domain Category
@@
 Entertainment Riddles, Jokes, Game Suggestions.
-Table 2: Details of Alpaca-Ti.
 Alpaca-Ti encompasses a diverse array of topics and also has a longer length and better instruction fine-tuning performance, designed to enhance LLM’s generalization capabilities.
 It has a total of 42,676 instances.
-One data sample is shown in Appendix Figure 9.
+One data sample is shown in Fig.
+4.
+Figure 4: One sample from Alpaca-Ti and ’-’ means nothing was entered.
 3.1.3 Safety-Prompts-Ti
-Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table 3, and the data size is 38,156.
-LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts Chao et al.
-(2024); Zheng et al.
-(2024); Sun et al.
-(2023).
-Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a dataset to simulate the process in advance.
-One data sample is shown in Appendix Figure 10.
+Safety-Prompts-Ti aims to improve the safety of large models on two major categories: Instruction Attacks and Typical Safety Scenarios, as shown in Table III, and the data size is 38,156.
+LLM’s representation space, harmful and harmless queries can be largely distinguished, but this is not noticeably enhanced by safety prompts [22, 23, 24].
 Instruction Attacks 15k
@@
 Others 4.7K
-Table 3: Statistical details of Instruction Attacks and Typical Safety Scenarios.
+TABLE III: Statistical Details of Instruction Attacks and Typical Safety Scenarios
+Human evaluation is a valuable approach; however, due to the extensive workload it entails, we opted to create a benchmark to simulate the process in advance. One data sample is shown in Fig. 5.
+Figure 5: One sample from Safety-Prompts-Ti.
 3.1.4 CValues-Ti
-CValues-Ti is the dataset, used to improve the security of Tibetan language models.
-Considering work Moore et al.
-(2024); Rozen et al.
-(2024); Xu et al.
-(2023), we propose this branch of a small dataset because LLM can pose risks or has negative social impacts.
+CValues-Ti is the benchmark, used to improve the security of Tibetan language models.
+Considering work [25, 27, 28], we propose this branch of a small benchmark because LLM can pose risks or has negative social impacts.
 For CValues-Ti, it has three categories (safe & reponsibility = rejection & positive advice; safe = rejection; unsafe = risk response).
 Then, under the same prompt, various types of responses can be categorized into positive and negative samples, differentiated by their levels of difficulty, so that there are three common combinations (safe + unsafe; safe & reponsibility + unsafe; safe& reponsibility + safe).
-Unlike the work Xu et al.
-(2023), We streamlined and condensed the instructions, and compared to their dataset (20K), our sub-dataset, CValues-Ti (2K), achieved equivalent results in subsequent experiments.
-One data sample is shown in Appendix Figure 11.
+Unlike the work [28], We streamlined and condensed the instructions, and compared to their benchmark (20K), our sub-benchmark, CValues-Ti (2K), achieved equivalent results in subsequent experiments.
+One data sample is shown in Fig.
+6.
+Figure 6: One sample from CValues-Ti.
 3.1.5 hh-rlhf-Ti
-hh-rlhf-Ti is also the first dataset that improves the model effect by introducing human intervention.
-It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF Bai et al.
-(2022); Kirk et al.
-, but we reset and summarize the previous dataset based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives.
-This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan.
-It has a total of 47,411 instances, and one data sample is shown in Appendix Figure 12.
-3.2 Sun-Shine
 Figure 7: The training process of the Sun-Shine, consists of three key stages: Pre-training, SFT, and DPO.
@@
 DPO adjusts the Sun-Shine to prioritize user preferences, resulting in outputs that are more aligned and appealing.
-Sun-Shine is built on the LLaMA 3 Touvron et al.
-(2023); Dubey et al.
-(2024) architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community.
-For more accessible training, we adopt the LLaMA 3.1 (8B) model Dubey et al.
-(2024).
-Additionally, we integrate NLP model translation with human annotation Üstün et al.
-(2024) as supplementary data to enhance Sun-Shine’s accuracy.
-As illustrated in Figure 1, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO).
-For training details, Appendix Table 6 has shown parameters’ setting.
-Notably, we adopted LoRA-based parameter-efficient fine-tuning Hu et al.
-for both SFT and DPO, which significantly reduced memory usage without sacrificing performance.
-3.2.1 Pre-Training
-During the Pre-training stage, the Sun-Shine is pre-trained on Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction.
-And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text.
-For details,
-3.2.2 Supervised Fine-Tuning
-Following pretraining, the model undergoes supervised fine-tuning Yu et al.
-(2021) with Alpaca-Ti and Safety-Prompts-Ti.
-During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
-3.2.3 Direct Preference Optimization
-In the stage of DPO Rafailov et al.
-(2024); Chao et al.
-(2024); Zheng et al.
-(2024), the model’s alignment and output quality are further optimized using preference data.
-The input data comes from CValues-Ti and hh-rlhf-Ti.
-Preference Data is generated, representing user preferences for certain types of outputs.
-The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
-4 Evaluation: TLUE Benchmark
-Unfortunately, existing benchmarks Li et al.
-(2024); Zhang et al.
-(2024); Sun et al.
-(2023) are available only in Chinese or English, with no purely Tibetan versions.
-To address this issue, we first translated the data and manually corrected them.
-This has been made into a new benchmark, TLUE Gao et al.
-(2025).
-Fortunately, our partnerships include Tibetan institutions based in Tibet.
-Unlike modern Tibetan, Our translation conversion strives to preserve the ancient grammar and unique linguistic features of traditional Tibetan culture.
-All data have been carefully validated by Tibet experts and reviewed by relevant institutions.
-TLUE has two branches: Ti-MMLU and Ti-Safetybench.
-In this paper, we use both of them to assess Sun-Shine’s Tibetan language ability.
-4.1 Ti-MMLU
-CMMLU Li et al.
-(2024) is a comprehensive Chinese benchmark that spans multiple disciplines, including natural sciences, social sciences, engineering, and the humanities, which consists of 11,528 questions distributed across 67 subjects, with each subject containing a minimum of 105 questions.
-These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions.
-The benchmark encompasses 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains.
-Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere.
-The evaluation matrix includes Response Rate, Accuracy, and Conditional Accuracy.
-Response rate measures the proportion of instances where LLMs provide a valid response.
-Accuracy indicates how often LLMs’ selected answer matches the correct one, while conditional accuracy assesses LLMs’ accuracy only when a valid response is given.
-4.2 Ti-Safetybench
-SafetyBench Zhang et al.
-(2024); Sun et al.
-(2023) comprises 11,435 diverse multiple choice questions spanning across 7 distinct categories of safety concerns: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality and Privacy and Property.
-Its sources are diverse, comprising questions from existing Chinese and English datasets, safety-related examinations, and ChatGPT-augmented samples.
-Although our prompts ask for direct answers, many models automatically analyze each option.
-To calculate these metrics, we use two methods: Direct Answer Calculation (DA) and Concern All Answer Calculation (CAA).
-DA evaluates the model’s ability to respond to Tibetan instructions by checking if its selected answer is valid and correct.
-CAA assesses the performance ceiling by filtering valid choices, removing full answer combinations, and verifying correctness based on the remaining choice.
-5 Experimental Result
-We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) Achiam et al.
-(2023), Llama-3.18B Dubey et al.
-(2024); Touvron et al.
-(2023), Qwen-2.5 (72B) Yang et al.
-(2024a) and DeepSeek-R1 Bi et al.
-(2024); Guo et al.
-(2025).
-Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis.
-5.1 Insights from Performance Results
-Ti-MMLU: as shown in Figure 2, Figure 3 and Appendix Table 7, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs .
-(a) Response Rate
-(b) Conditional Accuracy
-(c) Accuracy
-Figure 2: Ti-MMLU benchmark comparison: Sun-Shine vs. other LLMs.
-(a) The highest total number of indicators.
-(b) The second total number of indicators.
-Figure 3: Statistics of Sun-Shine ranks first and second based on the number of indicators, across all categories of the Ti-MMLU benchmark.
-For Response Rate, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1 (see Appendix Table 8 and Figure 3).
-Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training dataset.
-For Conditional Accuracy, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead (Appendix Table 9).
-Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities.
-Sun-Shine excels among most LLMs but lags slightly behind DeepSeek-R1 due to the larger dataset used to train the latter’s base model.
-For Accuracy, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
-This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient.
-Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
-Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance.
-Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine.
-Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
-Ti-Safetybench: as shown in Figure 4 and Appendix Table 11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods.
-Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA.
-GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
-(a) Comparison of LLMs Across DA
-(b) Comparison of LLMs Across CCA
-Figure 4: Ti-SafetyBench comparison: Sun-Shine vs. other LLMs.
-Overall, Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1.
-While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance.
-This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
-5.2 Sun-Shine for Tibetan Culture
-As shown in Appendix Figure 13, Appendix Figure 14, and Appendix Figure 15, Sun-Shine maintains a neutral tone when addressing common questions, avoiding exaggerated praise or references to Tibetan classics.
-In contrast, GPT-4o adopts a more amplified and majestic style, reminiscent of divine scripture, while DeepSeek-R1’s responses frequently allude to the bodhisattva Mañjuśrī, appearing slightly ”overfitted” with more refined yet noticeably embellished language than GPT-4o.
-Figure 5: Experimental results of classical Chinese’s output from Sun-Shine, GPT-4o and DeepSeek-R1.
-As shown in Figure 5, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese.
-This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records.
-The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
-Overall, Sun-Shine surpasses DeepSeek-R1 (671B) and GPT-4o in most performance indicators, with its classical Tibetan style attributed to the curated TIB-STC and training strategy, where human annotation and correction played a key role.
-5.3 Downstream Task
-Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan).
-BELU Papineni et al.
-(2002) and ROUGE Lin (2004) are used to evaluate this task.
-As shown in Table 4, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower.
-Table 5 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
-LLM B-4 R-1 R-2 R-l
-Sun-Shine 42.16 65.78 48.91 51.58
-GPT-4o 38.85 61.12 44.33 47.02
-DeepSeek-R1 41.85 65.49 47.73 51.04
-Qwen-2.5 35.20 58.59 41.25 44.73
-Table 4: Comparison of translation task indicators with other LLMs.
-Tibetan Language Proficiency Test: we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
-LLM A B C D
-Sun-Shine 63 61 44 42
-GPT-4o 45 47 53 51
-DeepSeek-R1 59 56 67 66
-Qwen-2.5 41 43 53 54
-Table 5: Comparison of Tibetan Language Proficiency Test with other LLMs.
-As shown in Table 5, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics.
-This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our dataset, TIB-STC, which includes a richer collection of ancient Tibetan texts.
-6 Conclusion
-In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC dataset, marking significant progress in Tibetan language processing.
-Sun-Shine, trained on the comprehensive TIB-STC dataset, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios.
-TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data.
-Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan.
-Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our dataset, TIB-STC, in providing a high-quality ancient Tibetan corpus.
-This work advances low-resource language models and underscores the importance of culturally specific datasets.
-Acknowledgments
-This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.
-Appendix A Appendix A
-A.1 Tibetan Components and their Composition
-Figure 6: This is the native tibetan syllable structure, which means raise or lift.
-It consists of 7 parts: prefix letter, root letter, superscript letter, subscript letter, vowel letter, suffix letter, and farther suffix letter.
-Figure 7: Characters belong to each component.
-A.2 Details of TIB-STC
-A.2.1 Corpus Subset
-Figure 8: The two data samples from the corpus of TIB-STC are from the categories of ”Buddhism” and ”Culture” respectively.
-A.2.2 Fine-Tuning Subset: Alpaca-Ti
-Figure 9: One sample from Alpaca-Ti and ’-’ means nothing was entered.
-A.2.3 Fine-Tuning Subset: Safety-Prompts-Ti
-Figure 10: One sample from Safety-Prompts-Ti.
-A.2.4 Alignment Subset: CValues-Ti
-Figure 11: One sample from CValues-Ti.
-A.2.5 Alignment Subset: hh-rlhf-Ti
+hh-rlhf-Ti is also the first benchmark that improves the model effect by introducing human intervention.
+It consists of two parts: ”harmless (23,822)” and ”helpful (23,589)” just like RLHF [29, 30], but we reset and summarize the previous benchmark based on the characteristics of Tibetan grammar, and introduce translation, which can either be integrated with or conducted prior to training for other objectives.
+This approach could be important in the future to prevent the development of intermediate, unaligned AI systems for Tibetan.
+It has a total of 47,411 instances, and one data sample is shown in Fig.
+8.
 Figure 8: One sample from hh-rlhf-Ti and ’…’ means content where a person interacts with an assistant that the assistant represents the LLM.
 During this process, people can change the content of the questions or the way of guidance.
-A.3 Hyperparameter Settings for Each Stage of Sun-Shine Training
+3.2 Sun-Shine
+Sun-Shine is built on the LLaMA-3 [2, 20] architecture, a collection of multilingual base models with 7 to 65 billion parameters, openly available to the research community.
+For more accessible training, we adopt the LLaMA 3.1 (8B) model [20].
+Additionally, we integrate NLP model translation with human annotation [21] as supplementary data to enhance Sun-Shine’s accuracy.
+As illustrated in Fig.
+7, the training process of Sun-Shine consists of three stages: Pre-training, Supervised Fine-Tuning (SFT), and Direct Preference Optimization (DPO).
+For training details, Table IV has shown parameters’ setting.
+Notably, we adopted LoRA-based parameter-efficient fine-tuning [15] for both SFT and DPO, which significantly reduced memory usage without sacrificing performance.
+Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages.
+LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO.
+All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1.
+The pretraining phase involved over 11B tokens with three full epochs.
+This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
 Parameter Training Stage
@@
 pref_loss Sigmoid
-Table 6: Training hyperparameters used in each stage of Sun-Shine’s development.
-Training was performed using bfloat16 precision and a batch size of 4096 tokens across all stages.
-LoRA adapters (rank=32) were applied to both attention and feed-forward layers during SFT and DPO.
-All stages used AdamW optimizer with a weight decay of 0.1 and a dropout rate of 0.1.
-The pretraining phase involved over 11B tokens with three full epochs.
-This configuration ensures both computational efficiency and stable convergence for low-resource language modeling.
-Appendix B Appendix
-B.1 Details of Experimental Results on TLUE Benchmark
-Note: the highest-ranking; the second-highest) (BLEU-4 = B-4; ROUGE-1 = R-1; ROUGE-2 = R-2; ROUGE-l = R-l) S-S = Sun-Shine; G-4o = GPT-4o; L-318B = Llama-3.18B; G-35T = GPT-3.5-Turbo; Q-2.5 = Qwen-2.5; Go1-m = GPT-o1-mini; DK-r1 = DeepSeek-r1; Offensiveness = OFF, Unfairness and Bias = UB, Physical Health = PH, Mental Health = MH, Illegal Activities = IA, Ethics and Morality = EM and Privacy and Property = PP.
-B.1.1 Overview of LLMs Experiments on the Branch of TLUE: Ti-MMLU
-Model Response Rate
-China Specific Social Sciences Humanities STEM Other
-Sun-Shine 0.5507 0.5588 0.5377 0.5481 0.5498
-GPT-4o 0.4932 0.5133 0.5107 0.4915 0.4957
-Llama-3.18B 0.1907 0.1998 0.2056 0.2082 0.2104
-GPT-3.5-turbo 0.0687 0.0741 0.0748 0.0806 0.0801
-Qwen-2.5 0.2563 0.2703 0.2736 0.2441 0.2419
-GPT-o1-mini 0.2050 0.2340 0.2333 0.2231 0.2241
-DeepSeek-R1 0.4206 0.4069 0.3884 0.3580 0.3482
-Model Conditional Accuracy
-China Specific Social Sciences Humanities STEM Other
-Sun-Shine 0.3208 0.3367 0.3477 0.3457 0.3489
-GPT-4o 0.2845 0.2871 0.2846 0.2746 0.2775
-Llama-3.18B 0.2510 0.2704 0.2662 0.2620 0.2597
-GPT-3.5-turbo 0.2489 0.2605 0.2656 0.2710 0.2731
-Qwen-2.5 0.2656 0.2655 0.2710 0.2866 0.2900
-GPT-o1-mini 0.2636 0.2700 0.2716 0.2742 0.2728
-DeepSeek-R1 0.3209 0.3482 0.3524 0.3584 0.3573
-Model Accuracy
-China Specific Social Sciences Humanities STEM Other
-Sun-Shine 0.1888 0.1742 0.1683 0.1614 0.1618
-GPT-4o 0.1594 0.1676 0.1634 0.1515 0.1535
-Llama-3.18B 0.0656 0.0628 0.0723 0.0662 0.0663
-GPT-3.5-turbo 0.0182 0.0201 0.0205 0.0277 0.0265
-Qwen-2.5 0.0691 0.0723 0.0744 0.0702 0.0709
-GPT-o1-mini 0.0475 0.0584 0.0598 0.0603 0.0606
-DeepSeek-R1 0.1863 0.1819 0.1705 0.1562 0.1534
-Table 7: Benchmarking Sun-Shine Against Leading LLMs on Ti-MMLU .
-For more detailed information about these three indicators, please refer to Appendix Table 8 (Response Rate), Appendix Table 9 (Conditional Accuracy) and Appendix Table 10 (Accuracy).
-The Category column lists the categories in order from top to bottom: China Specific, Social Sciences, Humanities, STEM, and Other.
-B.1.2 Details of Response Rate
+TABLE IV: Hyperparameters in Each Stage of Sun-Shine
+3.2.1 Pre-Training
+During the Pre-training stage, the Sun-Shine is pre-trained on Corpus, which helps the Sun-Shine learn general language representation and generation capabilities using a standard objective, such as Next Token Prediction.
+And the input consists of text data, and the goal is to enable Sun-Shine to understand and generate Tibetan text.
+For details,
+3.2.2 Supervised Fine-Tuning
+Following pretraining, the model undergoes supervised fine-tuning [18] with Alpaca-Ti and Safety-Prompts-Ti.
+During this phase, the model is fine-tuned to generate outputs based on specific instructions and inputs, whose objective is to align the model’s behavior with specific tasks or domains while ensuring safety and professional.
+3.2.3 Direct Preference Optimization
+In the stage of DPO [17, 22, 23], the model’s alignment and output quality are further optimized using preference data.
+The input data comes from CValues-Ti and hh-rlhf-Ti.
+Preference Data is generated, representing user preferences for certain types of outputs.
+The optimization process directly maximizes the likelihood of preferred outputs, aligning Sun-Shine more closely with user expectations and preferences.
+4 Evaluation: TLUE Benchmark
+Unfortunately, existing benchmarks like CMMLU [35] and Safetybench [36, 24] are available only in Chinese or English, with no purely Tibetan versions.
+To address this issue, we used TLUE [42], the first Tibetan language benchmark that includes Ti-MMLU and Ti-Safetybench.
+Their data were translated and manually corrected, preserving the nuances of traditional Tibetan culture.
+4.1 Ti-MMLU
+CMMLU [35] is a well-established Chinese benchmark that covers multiple disciplines, including natural sciences, social sciences, engineering, and the humanities.
+To adapt this benchmark for Tibetan, TLUE [42] leveraged an extended version, Ti-MMLU, where the original questions were translated into Tibetan and manually validated to preserve linguistic accuracy and cultural relevance.
+It consists of 11,528 questions spanning 67 subjects, with each subject containing at least 105 questions.
+These are divided into a few-shot development set of 5 questions and a test set comprising over 100 questions.
+The benchmark includes 17 STEM tasks, 13 humanities tasks, 22 social science tasks, and 15 tasks from other domains.
+Notably, 16 tasks are China-specific, meaning they either lack equivalents in other regions or may have contextually different answers elsewhere.
+The evaluation framework follows the original CMMLU metrics, including Response Rate, Accuracy, and Conditional Accuracy.
+Response rate measures the proportion of instances where LLMs provide a valid response.
+Accuracy evaluates the percentage of correct answers, while conditional accuracy assesses performance only when a valid response is given.
+4.2 Ti-Safetybench
+SafetyBench [36, 24] is a widely used benchmark that contains 11,435 multiple-choice questions across 7 distinct safety categories: Offensiveness, Unfairness and Bias, Physical Health, Mental Health, Illegal Activities, Ethics and Morality, and Privacy and Property.
+Its questions are sourced from diverse benchmarks in Chinese and English, supplemented with safety-related examination content and ChatGPT-augmented samples.
+To evaluate safety-related capabilities for Tibetan, TLUE [42] employed Ti-Safetybench, a Tibetan adaptation of SafetyBench.
+The original data were translated and manually validated by experts to ensure consistency with traditional Tibetan linguistic and cultural standards.
+The evaluation methodology incorporates two scoring methods: Direct Answer Calculation and Concern All Answer Calculation.
+DA assesses the model’s ability to provide valid and accurate responses, while CAA calculates the performance ceiling by filtering valid choices and verifying correctness after removing irrelevant combinations.
+Both methods ensure a comprehensive assessment of the model’s safety performance in Tibetan contexts.
+Note: as shown in Table V, in order to make the paper more aesthetically pleasing and easier for researchers to read, some specialized terms are abbreviated.
+Category Full Name Abbreviation
+STEM STEM
+Humanities Human
+Ti-MMLU Social Sciences Social
+Other Other
+China Specific China
+Offensiveness OFF
+Unfairness and Bias UB
+Physical Health PH
+Ti-Safetybench Mental Health MH
+Illegal Activities IA
+Ethics and Morality EM
+Privacy and Property PP
+Response Rate RR
+Accuracy ACC
+Evaluation Metrics Conditional Accuracy CA
+Direct Answer Calculation DA
+Concern All Answer Calculation CAA
+TABLE V: List of Abbreviations for Professional Terms
+5 Experimental Result
+We compare Sun-Shine with currently popular LLMs in GPT series (4o, o1-mini and 3.5-turbo) [4], Llama-3.18B [20, 2], Qwen-2.5 (72B) [34] and DeepSeek-R1 [37, 38].
+Additionally, we provide extensive details of the experiment, including the experimental indicators for each branch, along with a thorough analysis.
+As shown in Table VI, regarding open-source model execution, the Qwen-2.5 and DeepSeek-r1 were evaluated via their respective official APIs provided by the model developers.
+The LlaMA-3.18B was accessed and run using the LlaMA-API platform.
+These details will also be clearly documented in the final version to enhance transparency and reproducibility.
+LLM Temperature Top_p Stream
+GPT-3.5-turbo 1.0 1.0 False
+GPT-4o 1.0 1.0 False
+GPT-o1-mini 1.0 1.0 True
+DeepSeek-R1 1.0 None Rtue
+LlaMA-3.18B 0.6 0.9 False
+Qwen-2.5 0.7 0.8 False
+TABLE VI: Hyperparameters of LLM
+Also, as shown in Table VII, LLM and some indicators of downstream tasks, such as BLUE [40] and ROUGH [41], used to evaluate translation tasks, also uses abbreviations in some tables for easier reading.
+For the experimental metrics, the best metric is bolded and the suboptimal metric is underlined.
+Category Full Name Abbreviation
+LLM GPT-3.5-turbo G-35T
+GPT-4o G-4o
+GPT-o1-mini Go1-m
+DeepSeek-R1 DK-r1
+LlaMA-3.18B L-318B
+Qwen-2.5 Q-2.5
+Sun-Shine S-S
+Evaluation Metrics BLEU-4 B-4
+ROUGE-1 R-1
+ROUGE-2 R-2
+ROUGE-l R-l
+TABLE VII: List of Abbreviations for Professional Terms
+(a) Response Rate
+(b) Conditional Accuracy
+(c) Accuracy
+Figure 9: Ti-MMLU benchmark comparison: Sun-Shine vs. other LLMs.
+5.1 Insights from Performance Results
+Ti-MMLU: as shown in Fig. 9, Sun-Shine, GPT-4o and DeepSeek-R1 outperform all other LLMs.
+For Response Rate, as shown in Table VIII, Sun-Shine and GPT-4o significantly outperform all other LLMs, including DeepSeek-R1.
+Sun-Shine leads in every sub-branch, demonstrating superior responsiveness to Tibetan instructions due to its extensive training benchmark.
 Category Response Rate
@@
 elementary_IT 0.5462 0.5378 0.2353 0.1050 0.1975 0.2353 0.2647
-Table 8: Response Rate Comparison: Sun-Shine vs. Other LLMs.
-B.1.3 Details of Conditional Accuracy
+TABLE VIII: Response Rate: Sun-Shine vs. Other LLMs
+For Conditional Accuracy, as shown in Table IX, Sun-Shine and DeepSeek-R1 lead, with DeepSeek-R1 slightly ahead.
+Although neither ranks first in every indicator, both maintain consistently high scores, reflecting their strong knowledge capabilities.
 Category Conditional Accuracy
@@
 elementary_IT 0.3185 0.3906 0.2500 0.3600 0.2128 0.2874 0.3452
-Table 9: Conditional Accuracy Comparison: Sun-Shine vs. Other LLMs.
-B.1.4 Details of Accuracy
+TABLE IX: Conditional Accuracy: Sun-Shine vs. Other LLMs
+For Accuracy, as shown in Table X, Sun-Shine outperforms DeepSeek-R1 in China-Specific, STEM, and Other categories, while DeepSeek-R1 leads in the remaining areas.
 Category Accuracy
@@
 elementary_IT 0.1092 0.2102 0.0588 0.0378 0.0420 0.0798 0.1387
-Table 10: Accuracy Comparison: Sun-Shine vs. Other LLMs.
-B.2 Overview of LLMs Experiments on the Branch of TLUE: Ti-Safetybench
-Model Category
+TABLE X: Accuracy: Sun-Shine vs. Other LLMs.
+This highlights Sun-Shine’s superior Tibetan processing ability, influenced by the data distribution in TIB-STC, underscoring that model architecture alone is insufficient.
+Effective data utilization and tailored training strategies enable Sun-Shine (8B) to surpass DeepSeek-R1 (671B, MOE LLM).
+(a) The highest total number of indicators
+(b) The second total number of indicators
+Figure 10: Statistics of Sun-Shine ranks first and second based on the number of indicators, across all categories of the Ti-MMLU benchmark.
+As shown in Fig.
+10, we also compiled statistics on various indicators of Sun-Shine in each sub-branch.
+Whether it is the best or the second best indicator, Sun-Shine maintains a large advantage in total, whether compared with DeepSeek-R1 or Chat-4o.
+Especially in the sub-category, “Other”, Sun-Shine obtained the best indicators of all accuracy.
+The remaining sub-categories also outperform DeepSeek-r1 overall.
+Only in the indicator ”Conditional Accuracy”, Sun-Shine is slightly lower than the latest DeepSeek-R1.
+Also, as shown in Table XI, we also calculated the average of different metrics for different large models on different categories.
+Sun-Shine is either first or second in all categories and all indicators.
+Among other LLMs, GPT-o1-mini and GPT-3.5-Turbo show weaker performance.
+Qwen-2.5 surpasses LLaMA-3.18B but falls short compared to GPT-4o, DeepSeek-R1, and Sun-Shine.
+Notably, GPT-o1-mini performs significantly worse than DeepSeek-R1 despite its 100-fold higher inference cost.
+LLM Response Rate
+China Social Human STEM Other
+Sun-Shine 0.5507 0.5588 0.5377 0.5481 0.5498
+GPT-3.5-turbo 0.0687 0.0741 0.0748 0.0806 0.0801
+GPT-4o 0.4932 0.5133 0.5107 0.4915 0.4957
+GPT-o1-mini 0.2050 0.2340 0.2333 0.2231 0.2241
+Llama-3.18B 0.1907 0.1998 0.2056 0.2082 0.2104
+Qwen-2.5 0.2563 0.2703 0.2736 0.2441 0.2419
+DeepSeek-R1 0.4206 0.4069 0.3884 0.3580 0.3482
+LLM Conditional Accuracy
+China Social Human STEM Other
+Sun-Shine 0.3208 0.3367 0.3477 0.3457 0.3489
+GPT-3.5-turbo 0.2489 0.2605 0.2656 0.2710 0.2731
+GPT-4o 0.2845 0.2871 0.2846 0.2746 0.2775
+GPT-o1-mini 0.2636 0.2700 0.2716 0.2742 0.2728
+Llama-3.18B 0.2510 0.2704 0.2662 0.2620 0.2597
+Qwen-2.5 0.2656 0.2655 0.2710 0.2866 0.2900
+DeepSeek-R1 0.3209 0.3482 0.3524 0.3584 0.3573
+LLM Accuracy
+China Social Human STEM Other
+Sun-Shine 0.1888 0.1742 0.1683 0.1614 0.1618
+GPT-3.5-turbo 0.0182 0.0201 0.0205 0.0277 0.0265
+GPT-4o 0.1594 0.1676 0.1634 0.1515 0.1535
+GPT-o1-mini 0.0475 0.0584 0.0598 0.0603 0.0606
+Llama-3.18B 0.0656 0.0628 0.0723 0.0662 0.0663
+Qwen-2.5 0.0691 0.0723 0.0744 0.0702 0.0709
+DeepSeek-R1 0.1863 0.1819 0.1705 0.1562 0.1534
+TABLE XI: Benchmarking Sun-Shine Against Leading LLMs on Ti-MMLU.
+Ti-Safetybench: as shown in Fig.
+11, Sun-Shine was evaluated against leading LLMs on Ti-SafetyBench across safety-related categories using DA and CCA filtering methods.
+Compared to GPT-4o, LLaMA-3.18B, GPT-3.5-Turbo, Qwen-2.5, GPT-o1-mini, and DeepSeek-R1, Sun-Shine excels in IA, OFF, UB, and MH categories, while DeepSeek-R1 performs well under CCA.
+GPT-4o and GPT-3.5-Turbo show moderate results, with LLaMA-3.18B and Qwen-2.5 scoring lower in safety metrics.
+(a) Comparison of LLMs Across DA
+(b) Comparison of LLMs Across CCA
+Figure 11: Ti-SafetyBench: Sun-Shine vs. other LLMs
+As shown in Table XII, it presents a comprehensive comparison of leading LLMs on the Ti-Safetybench, evaluating their performance across key safety-related categories.
+Notably, Sun-Shine outperforms all baseline models under both DA and CCA tasks.
+In the DA setting, Sun-Shine achieves the highest exact match (EM) score (0.292) and leads in IA, OFF, and UB, while maintaining a significantly higher filter rate (59.20%) than GPT-4o (38.59%) and others.
+In the CCA setting, Sun-Shine demonstrates even stronger results, ranking first across nearly all categories, including EM (0.483), IA (0.455), mental health (MH, 0.507), and UB (0.466).
+These results highlight Sun-Shine’s superior ability to generate safe, accurate, and instruction-aligned responses, setting a new benchmark in safety-aware language modeling.
+LLM Category
 Filter Invaild EM IA MH OFF PH PP UB
@@
 DeepSeek-R1 28.14% 0.558 0.457 0.500 0.429 0.511 0.339 0.457
-Table 11: Benchmarking Sun-Shine Against Leading LLMs on Ti-Safetybench.
-B.3 Output of Experimental Results from LLMs
-Figure 13: Example of Sun-Shine’s output.
-Figure 14: Example of GPT-4o’s output.
-Figure 15: Example of DeepSeek-R1’s output.
+TABLE XII: Benchmarking Sun-Shine Against Leading LLMs on Ti-Safetybench
+Overall: Sun-Shine demonstrates slightly stronger security capabilities in Tibetan language processing than DeepSeek-R1.
+While DeepSeek-R1’s base model surpasses LLaMA-3.18B in security, Sun-Shine’s extensive pre-training and fine-tuning, supported by the data advantage of TIB-STC, achieve comparable performance.
+This not only validates Sun-Shine’s capabilities but also underscores the strengths of TIB-STC.
+5.2 Sun-Shine for Tibetan Culture
+As shown in Fig.
+12, Sun-Shine significantly outperforms DeepSeek-R1 and GPT-4o in classical Chinese.
+This superiority stems from its training on TIB-STC, which includes ancient poems, essays, and historical records.
+The manual revision process preserved Tibetan grammar, sentence structure, and style, giving Sun-Shine’s output a distinctly classical character, unlike the modern Tibetan style used by DeepSeek-R1 and GPT-4o.
+Figure 12: Experimental results of classical Chinese’s output from Sun-Shine, GPT-4o and DeepSeek-R1.
+As shown in Fig.
+13(a), Fig.
+13(b), and Fig.
+13(c), we observe clear stylistic distinctions in how different LLMs respond to a culturally sensitive prompt, ”Please write a poem in praise of father”, in Tibetan, Chinese, and English.
+(a) Output of Sun-Shine
+(b) Output of GPT-4o
+(c) Output of DeepSeek-R1
+Figure 13: Ti-SafetyBench: Sun-Shine vs. other LLMs
+Sun-Shine generates outputs that are notably neutral, emotionally grounded, and culturally restrained across all three languages.
+The Tibetan response is respectful and sincere, avoiding ornate metaphors or religious references.
+In Chinese, the model articulates gratitude with clarity and warmth, emphasizing familial support without exaggeration.
+The English version maintains a reflective tone, expressing appreciation without invoking religious or hyperbolic imagery.
+This demonstrates Sun-Shine’s alignment with real-world communication norms and its ability to generalize across linguistic and cultural contexts.
+In contrast, GPT-4o exhibits a distinctly elevated and ceremonial tone.
+Its Tibetan and Chinese outputs are imbued with grandiloquent phrases and spiritual imagery.
+For example, the model refers to the father as a “beacon of happiness” and ”born on the edge of happiness,” which, while poetic, may not align with typical expressions of filial appreciation.
+The English version similarly includes metaphor-heavy language, such as “the victory banner is my prayer,” which resembles religious scripture more than natural expression.
+This suggests a stylistic bias toward exalted, sometimes deific, portrayals.
+DeepSeek-R1, on the other hand, appears to have an even stronger inclination toward religious and cultural iconography.
+It references “Mañjuśrī’s Father” and constructs its Tibetan output with vocabulary and structure closely mirroring classical Buddhist literature.
+The Chinese response is elegant yet heavily stylized, and the English output reads like a translated verse from scripture rather than a naturally composed poem.
+While aesthetically refined, this stylistic preference may reflect overfitting to religious texts in training data, limiting the model’s adaptability to general-purpose use.
+In summary, Sun-Shine’s responses are more balanced and less prone to embellishment, demonstrating cultural sensitivity and instruction alignment.
+GPT-4o and DeepSeek-R1, while fluent and creative, tend to overgenerate in stylistically loaded contexts, risking a mismatch with the intended pragmatic tone.
+This comparison underscores Sun-Shine’s strength in producing appropriate, grounded, and respectful language across multiple languages and cultural domains.
+5.3 Downstream Task
+Tibetan Language Proficiency Test: we connect these LLMs to this test, Tibet Autonomous Region Tibetan-Chinese Translation-2024111http://cn.zyw.xizang.gov.cn/ (100-point system).
+LLM A B C D
+Sun-Shine 63 61 44 42
+GPT-4o 45 47 53 51
+DeepSeek-R1 59 56 67 66
+Qwen-2.5 41 43 53 54
+TABLE XIII: Comparison of Tibetan Language Proficiency Test with other LLMs.
+As shown in Table XIII, Sun-Shine achieves its highest performance on the more challenging A and B exams, which contain a significant amount of ancient Tibetan content, while its performance is comparatively lower on C and D exams that feature more modern Tibetan content, including sensitive recent political topics.
+This score distribution highlights Sun-Shine’s strong understanding of ancient Tibetan and also demonstrates the superiority of our benchmark, TIB-STC, which includes a richer collection of ancient Tibetan texts.
+Translation: we performed a downstream task, Tibetan-Chinese translation (Chinese to Tibetan).
+BELU [40] and ROUGE [41] are used to evaluate this task.
+As shown in Table XIV, Sun-Shine achieves the highest scores across all metrics, demonstrating superior translation capabilities, followed by DeepSeek-R1, with GPT-4o and Qwen-2.5 scoring relatively lower.
+Table 12 can laso prove this point that the generated Tibetan is closer to the ancient style and conforms to the annotation.
+LLM B-4 R-1 R-2 R-l
+Sun-Shine 42.16 65.78 48.91 51.58
+GPT-4o 38.85 61.12 44.33 47.02
+DeepSeek-R1 41.85 65.49 47.73 51.04
+Qwen-2.5 35.20 58.59 41.25 44.73
+TABLE XIV: Comparison of Translation Task
+6 Conclusion
+In this paper, we present two pioneering contributions: the Sun-Shine large language model and the TIB-STC benchmark, marking significant progress in Tibetan language processing.
+Sun-Shine, trained on the comprehensive TIB-STC benchmark, excels in Tibetan language tasks, demonstrating expertise in Tibetan culture and strong performance in low-resource NLP scenarios.
+TIB-STC , the first large-scale Tibetan corpus, addresses data scarcity by encompassing literature, religious texts, news, and conversational data.
+Benchmark evaluations highlight Sun-Shine’s superiority over LLMs in accuracy, response rate, and cultural expression, especially in classical Tibetan.
+Leading performance in downstream tasks further demonstrates that Sun-Shine possesses the capabilities of a qualified ancient Tibetan expert, while also highlighting the superiority of our benchmark, TIB-STC, in providing a high-quality ancient Tibetan corpus.
+This work advances low-resource language models and underscores the importance of culturally specific benchmarks.
+Acknowledgment
+This work was supported in part by the National Science and Technology Major Project under Grant 2022ZD0116100, in part by the Tibetan Natural Science Foundation under Grant XZ202201ZR0054G.